OpenStudy (mtalhahassan2):

State where possible vector, parametric, and symmetric equations for each of the following lines. e. the line passing through the points X(-4,3,0) and O(0,0,0). f. the line passing through the point Q(1,2,4) and parallel to the z-axis.

OpenStudy (mtalhahassan2):

For e, XO= (0,0,0)-(-4,3,0) = (4,-3,0) Vector equation r= t(4,-3,0) parametric equation x=4t y=-3t z=0 Symetric Equation t=x/4 , t=y/-3 , z=0

OpenStudy (mtalhahassan2):

@ParthKohli

OpenStudy (mtalhahassan2):

@misty1212